# PyTomography: A Python Library for Medical Image Reconstruction

### v4

@@ Abstract
There is a need for open-source libraries in emission tomography that (i) use modern and popular backend code to encourage community contributions and (ii) offer support for the multitude of reconstruction algorithms available in recent literature, such as those that employ artificial intelligence. The purpose of this research was to create and evaluate a GPU-accelerated, open-source, and user-friendly image reconstruction library, designed to serve as a central platform for the development, validation, and deployment of various tomographic reconstruction algorithms. PyTomography was developed using Python and inherits the GPU-accelerated functionality of PyTorch and parallelproj for fast computations. Its flexible and modular design decouples system matrices, likelihoods, and reconstruction algorithms, simplifying the process of integrating new imaging modalities using various python tools. Example use cases demonstrate the software capabilities in parallel hole SPECT and listmode PET imaging. Overall, we have developed and publicly share PyTomography, a highly optimized and user-friendly software for medical image reconstruction, with a class hierarchy that fosters the development of novel imaging applications.
@@
C1 Current code version v3.0.0
@@
C7 Compilation requirements, operating environments & dependencies python >=>=> = 3.11
@@ 1 Motivation and significance
The unique purpose of our present research and effort is to build upon recent efforts in GPU-accelerated reconstruction [16, 17, 18, 19] and AI-based reconstruction [7, 8, 9] to establish an open source medical image reconstruction programming framework written in a popular programming language that facilitates collaboration. Valuable past efforts have focused on CUDA programming in languages such as C/C++ to accelerate reconstruction times in SPECT [16] and PET [16, 17] reconstruction through the use of multi-threading. While some contributors to an open-source framework may prefer to use C/C++ programming, others may be unfamiliar with these languages; any tool that prioritizes community contributions should permit both options. Some of the advantages of a python framework are that (i) it is a widely known programming language and thus encourages community involvement and (ii) it permits the embedding of C/C++ code for particular applications. Python can directly utilize CUDA acceleration via the PyTorch library, for example, allowing users unfamiliar with C/C++ code to achieve GPU acceleration.
@@
To this end, we developed the python library PyTomography with the priorities to: (i) implement standard and traditional imaging modalities and reconstruction algorithms, (ii) implement and disseminate of recent research developments, such as the deep image prior [8], and (iii) encourage community involvement via extensive documentation and user-friendly tutorials. Based on the motivations previously described, the backend of library has been developed using PyTorch because it maintains a similar functional structure to the widely used NumPy library. Other libraries are used for more particular tasks; the python library paralleproj [19], which was developed with CUDA C backend code, is used for line integral computation in PET projectors. In general, code written in any language can be integrated in PyTomography provided a python interface is available. While the present focus of PyTomography is the development and validation of reconstruction algorithms for SPECT and PET, implementation of other imaging modalities and reconstruction algorithms can be added using the building blocks provided.
@@
Overall, the aim of this study is to introduce the community to the PyTomography package, demonstrate its current capabilities and flexibility, and to encourage community involvement in the continuation of its development. To demonstrate the versatility of the software, example applications are explored in (i) SPECT reconstruction using PyTorch based rotate+sum projectors, and (ii) listmode (LM) time-of-flight (TOF) PET reconstruction. In what follows, we elaborate on our methods and results for PyTomography, which is made publicly available to download.
@@ 2.1 Software scope and functionality
A brief overview of the mathematical paradigm of medical imaging is presented in order to motivate the software architecture. The mathematical conventions used in this paper are as follows. Vectors are represented by lower case letters, while linear operators are represented by upper case letters. The product of two vectors $vw$ and division of two vectors $v/w$ are inferred to be point-wise operations. Components of a vector are given by $v_{i}$, and components of linear operators are given by $A_{ij}$. When $A1$ or $A^{T}1$ is written, “1” is to be interpreted as a vector (in the appropriate vector space) with all elements equal to 1.
@@ 2.1.1 Projectors
Linear imaging systems can be characterized by a system matrix $H:\mathcal{U}\to\mathcal{V}$. Its components $H_{ij}$ represent the contribution of element $j$ in object space to element $i$ in projection space. In emission tomography, analytical modeling of the data expectation $\bar{g}$ can be achieved via $\bar{g}=Hf+s$ where $s$ is an additive term containing estimates for scattered events (SPECT) or scatter and random coincidences (PET). In CT, analytical modeling is achieved via $\bar{g}=be^{-Hf}$ where $b$ denotes the rate of emissions from the transmission source to a particular measurement. The operation of $H$ on an object vector is known as forward projection (FP), while the operation of $H^{T}$ on a projection vector is known as back projection (BP). Example implementations available in PyTomography are listed below:
@@
    where $n$ designates the projection angle, $\mathcal{T}_{\text{rot}}^{(n)}$ implements a rotation, $\mathcal{T}_{\text{trans}}^{(n)}$ implements a translation that accounts for non-central projections, $\mathcal{T}_{\text{att}}^{(n)}$ implements the adjustment for photon attenuation, $\mathcal{T}_{\text{cdr}}^{(n)}$ implements depth-dependent collimator blurring, $\mathcal{P}_{x}$ is summation along the $+x$ axis, and $\mathcal{T}_{\text{fov}}^{(n)}$ applies a mask in projection space that accounts for a finite field of view, $v_{n}$ is a unit vector of length $N$ with 1 in component $n$, and $\otimes$ is the outer product.
@@
    where $\mathcal{T}_{rr}$ implements object-space resolution modeling (typically via a uniform Gaussian kernel), $\mathcal{P}_{\Sigma}$ implements line integral computation (listmode/sinogram), $\mathcal{T}_{\eta}$ adjusts for detector normalization, and $\mathcal{T}_{\omega}$ adjusts for attenuation effects.
PyTomography can be used to construct projectors that extend projection space; such functionality is necessary, for example, when modeling multiple photopeaks in SPECT or when using motion correction. In such situations, the extended projection space has dimensionality $(N,...)$ where ... is the original dimensionality of projection space, and $N$ represents the number of photopeaks or time gates during acquisition. Mathematically, these projectors can be written as
@@
    Multiple SPECT Photopeaks: In this scenario, $A_{n}$ are scaling factors that adjust for detector sensitivities and branching ratios, $B_{n}$ are identity mappings, and $H_{n}$ represents the system matrix corresponding to a particular photopeak. The extended projection space has dimensions $(N,...)$ where $N$ is the number of different photopeaks.
@@ 2.1.2 Likelihoods
A statistical likelihood function $\tilde{L}(g|\bar{g})$ is proportional to the probability of observing an actual realization of data $g$ given expectation $\bar{g}$. Since $\bar{g}$ is approximated using $f$ and $H$, it follows that $\tilde{L}(g|\bar{g})=\tilde{L}(g|f,H,...)$, where the ellipsis “...” contains additional parameters such as a scatter estimate $s$; the ellipsis will be assumed implicit for the remainder of this paper. Many iterative reconstruction algorithms seek to maximize the likelihood function via gradient ascent by exploiting the gradient $\nabla_{f}\tilde{L}(g|f,H)$. Since (i) computation of $\nabla_{f}\log\tilde{L}(g|f,H)$ is more tenable than $\nabla_{f}\tilde{L}(g|f,H)$ and (ii) maximization of $\log\tilde{L}(g|f,H)$ is equivalent to maximization of $\tilde{L}(g|f,H)$, the statistical “log-likelihood” is nearly always considered instead. In PyTomography, the “likelihood” $L(g|f,H)$ refers to an objective function to be maximized; as such, the “likelihood” $L(g|f,H)$ is often used interchangeably with the statistical “log-likelihood” $\log\tilde{L}(g|f,H)$ for a particular modality. Likelihood functions for different modalities are shown in Table 2.
Table 2: Likelihoods and corresponding gradients for various medical imaging modalities; $s$ represents a scatter estimate used in SPECT/PET while $b$ represents a scaling factor used in CT. The CT likelihood corresponds to an ideal scenario where the number of counts in each sinogram bin follows a Poisson distribution; this is not necessarily the case in real systems with complex electronic noise.
Likelihood $L(g|f,H)$ $\nabla_{f}L(g|f,H)$
SPECT/PET $\sum_{i}\left(g_{i}\log\left[Hf+s\right]_{i}-\left[Hf+s\right]_{i}\right)$ $H^{T}\left(\frac{g}{Hf+s}\right)-H^{T}1$
CT $\sum_{i}\left(g_{i}\log\left[be^{-Hf}\right]_{i}-\left[be^{-Hf}\right]_{i}\right)$ $H^{T}(be^{-Hf}-g)$
Weighted MSE $-\frac{1}{2}\sum_{i}\left[W\left(g-Hf\right)^{2}\right]_{i}$ $H^{T}W(g-Hf)$
@@ 2.1.3 Reconstruction Algorithms
where the ellipsis “...” includes all additional hyperparameters required for the algorithm; the ellipsis will be assumed implicit for the remainder of this paper. The main algorithms in PyTomography are preconditioned gradient ascent (PGA) algorithms; these algorithms aim to maximimize the likelihood via iterative updates that take the form
@@
where $n$ is the subiteration index (used as a superscript and not an exponent for all variables), $f^{n}$ is the object estimate of subiteration $n$, $g^{n}$ specifies a subset of data used in subiteration $n$, $H^{n}$ is a system matrix projecting only to the current subset, $C^{n}(f^{n})$ is a preconditioner factor that depends on the object estimate, $V(f^{n})$ is a prior function used to regularize the algorithm, and $\beta^{n}=c^{n}\beta$ is a scaling factor for the prior function (scaled by the fraction of subset data $c^{n}$). These algorithms partition the data $g$ into $M$ approximately equal subsets; a single iteration of an algorithm is defined as the number of subiterations where all subsets have been used once. Various PGA algorithms available in PyTomography are shown in Table 3; they are characterized by their preconditioner factor. Certain algorithm names in the literature correspond to specific realizations of the generic algorithms in Table 3 with particular likelihoods in Table 2. The OS-SART algorithm, for example, corresponds to the use of OSEM with the weighted MSE likelihood (with weights $W=1/(H1)$).
Table 3: Sample algorithms available in PyTomography: shown are ordered subset expectation maximum (OSEM), ordered subset maximum a posteriori one step late (OSMAPOSL), block sequential regualized expectation maximum (BSREM), and kernelized expectation maximum (KEM). $n$ is used as a superscript (not an exponent) for all variables. For BSREM, $\alpha^{n}$ is a relaxation parameter.
Algorithm Preconditioner factor $C^{n}(f^{n},H^{n})$
OSEM [20] $f^{n}/\left(\left(H^{n}\right)^{T}1\right)$ and $V(f^{n})=0$
OSMAPOSL [21] $f^{n}/\left(\left(H^{n}\right)^{T}1+\nabla_{f}V(f^{n})\right)$
BSREM [22] $f^{n}\alpha^{n}/\left(c^{n}H^{T}1\right)$
@@
where $w_{rs}$ provides a weighting factor between the locations voxel $r$ and voxel $s$ and $\phi(f_{r},f_{s})$ characterizes a weighting factor between the observed image values at voxel $r$ and voxel $s$. When $w_{rs}$ is non-zero only for neighbouring voxels, the prior is known as a nearest neighbour prior; various nearest neighbour priors implemented in PyTomography are shown in Table 4. While the weighting factor $w_{rs}$ is typically proportional to inverse Euclidean distance, it is also possible to allow it to depend on an additional anatomical image $\chi$ whereby $w_{rs}=w_{rs}(\chi_{r},\chi_{s})$ Since anatomical information is used, this is commonly referred to as an anatomical prior (AP); variations of this technique have recently become popular in clinical practice due to enhanced lesion quantitation and detectability in bone SPECT/CT [6]. In this paper, use of an AP refers to using only the 8 most similar neighbours based on absolute differences of HU values in a CT scan or attenuation map.
Table 4: Sample nearest neighbour functions available in PyTomography; these priors are characterized by their image weighting function $\phi(f_{r},f_{s})$.
Prior $\phi(f_{r},f_{s})$
Quadratic $\left((f_{r}-f_{s})/\delta\right)^{2}/4$
Log-cosh $\tanh((f_{r}-f_{s})/\delta)$
Relative difference penalty (RDP) [23] $\left((f_{r}-f_{s})^{2}\right)/\left(f_{r}+f_{s}+\gamma|f_{r}-f_{s}|\right)$
@@ 2.2 Software architecture
    Data/Metadata. The standard data type used to represent images and raw projection data is the tensor class of PyTorch; data can be stored on CPU or GPU. Metadata classes contain all auxiliary information needed for image reconstruction pertaining to a particular imaging system, such as voxel size, spatial dimensions, and information pertaining to the raw data of a particular modality.
@@
    Projectors. Projectors are mathematical operators that map between object and projection space. The system matrix is an example of a projector that models a particular imaging system. While the input/output datatype to projectors is the PyTorch tensor class, internal code can utilize functionality from other libraries. Projector subclasses inherit from a generic “SystemMatrix“ class and implement a “forward”, “backward”, and “compute_normalization_factor” method. Projectors used in subset based algorithms should additionally implement “set_n_subsets”, “get_projection_subset” and ”get_weighting_subset”; the second method takes in projection data and returns subset data based on how the projector defines subsets, and the third gets the corresponding weight ($c^{n}$) for scaling preconditioners.
@@
    Likelihoods. Likelihoods specify the likelihood function that characterizes the acquired data. Likelihoods $L(g|H,f)$ are initialized using a system matrix $H$, projection data $g$, and any additional dependencies; once initialized they act as functions of an image estimate $f$. Likelihoods inherit from a generic “Likelihood” class and should implement the “compute_gradient” method. Higher order derivatives of the likelihood may also be defined as seperate methods within the class.
@@
    Algorithms. Algorithms are used for image reconstruction, and are designed to interface with all imaging modalities. Statistical iterative algorithms require as input a likelihood instance, and may take as optional arguments prior functions and callbacks. Unlike projectors and likelihoods, algorithms do not inherit from a generic base class. In particular, however, linear PGAAs inherit from a generic “LinearPreconditionedGradientAscentAlgorithm” class and simply need to implement the “_linear_preconditioner_factor” and optionally reimplement the “__init__” method to take in required arguments.
PyTomography also has an input/output (I/O) module for reading various file formats; it currently supports SIMIND, DICOM (SPECT/CT), GATE (PET), and HDF5 listmode (PET) formats. Data reading consists of loading the generated files and constructing the appropriate data/metadata instances within the PyTomography framework. The I/O module is also used to read internal files of the library, such as coefficients to convert Hounsfield Units to attenuation coefficients, files containing SPECT collimator information, and files containing PET scanner geometry.
Novel development of a projector with associated data sources should thus consist of the following (i) implementation of metadata classes, (ii) I/O code to load data/metadata from the data sources, and (iii) development of a projector to interface with the data/metadata. Metadata class instances may be initialized by various data sources; for example, the same “SPECTProjMeta” class is initialized from SIMIND or DICOM I/O data loading functions, even though they represent different data sources. Metadata is thus linked to a modality and not to a specific data source.
Deployment of the software was designed to maximize accessibility. In addition to placing the backend code on GitHub, the library was also uploaded to the Python Package Index (PyPi) so users can easily install the library using the simple “pip install pytomography” terminal command. A readthedocs documentation website was created that also contains a comprehensive set of tutorials in jupyter notebook format. While many of the tutorials are aimed at training end users, there also exist developer tutorials for implementing new projectors in the software. A discourse website was developed for two purposes: (i) to allow end users to ask questions about how to use the software and (ii) to permanently store and make public the questions/answers as they may be useful for other users of the library.
Figure 1: Basic overview of the workflow of PyTomography when using statistical algorithms (left) with code representing a specific SPECT reconstruction use case (right). Shown are the minimum required inputs to the “__init__” function. the minimum required methods for each class, and sample subclasses available in the library. The “…” represents additional arguments that may be present in subclasses. In the SPECT use case, for example, the additional input argument “obj2obj_transforms” is implemented to take in operations that implement attenuation and PSF modeling. Not shown in the sample code: importing required libraries, loading the photopeak/scatter data, creating the attenuation/PSF transforms, and specifying the filepath “file_NM”.

### v5

@@ Abstract
There is a need for open-source libraries in emission tomography that (i) use modern and popular backend code to encourage community contributions and (ii) offer support for the multitude of reconstruction techniques available in recent literature, such as those that employ artificial intelligence. The purpose of this research was to create and evaluate a GPU-accelerated, open-source, and user-friendly image reconstruction library, designed to serve as a central platform for the development, validation, and deployment of various tomographic reconstruction algorithms. PyTomography was developed using Python and inherits the GPU-accelerated functionality of PyTorch and parallelproj for fast computations. Its flexible and modular design decouples system matrices, likelihoods, and reconstruction algorithms, simplifying the process of integrating new imaging modalities using various python tools. Example use cases demonstrate the software capabilities in parallel hole SPECT and listmode PET imaging. Overall, we have developed and publicly share PyTomography, a highly optimized and user-friendly software for medical image reconstruction, with a class hierarchy that fosters the development of novel imaging applications.
@@
C1 Current code version v3.2.3
@@
[…]
@@ 1 Motivation and significance
The unique purpose of our present research and effort is to build upon recent efforts in GPU-accelerated reconstruction [16, 17, 18, 19] and AI-based reconstruction [7, 8, 9] to establish an open source medical image reconstruction programming framework written in a popular programming language that facilitates collaboration. Valuable past efforts have focused on CUDA programming in languages such as C/C++ to accelerate reconstruction times in SPECT [16] and PET [16, 17] reconstruction through the use of multi-threading. While some contributors to an open-source framework may be comfortable with C/C++ programming, others may be unfamiliar with these languages; any tool that prioritizes community contributions should be designed with a front end application programming interface (API) that permits either of these options. Some of the advantages of a python front end API are that (i) it is a widely known programming language and thus facilitates community use, (ii) it enables direct interaction with other python libraries such as PyTorch for deep learning, and (iii) it enables backend development and contributions from other programming languages if necessary. This may address limitations of the C/C++ front end libraries STIR and CASToR, which are currently the most popular open-source reconstruction frameworks in emission tomography.
@@
To this end, we developed the python library PyTomography with the priorities to: (i) implement standard and traditional imaging modalities and reconstruction algorithms, (ii) disseminate recent research developments, such as the deep image prior [8], and (iii) encourage community involvement via extensive documentation and user-friendly tutorials. Based on the motivations previously described, the main functionality of the library has been developed using PyTorch. Other libraries are used for more particular tasks; the python library paralleproj [19] is used for line integral computation in PET projectors. While the present focus of PyTomography is the development and validation of reconstruction algorithms for SPECT and PET, implementation of other imaging modalities and reconstruction algorithms can be added using the building blocks provided.
In addition to the examples provided in this paper, PyTomography has support for high energy SPECT reconstruction via interface with the SPECTPSFToolbox [20], reconstruction using multiple SPECT photopeaks simultaneously, and region-based uncertainty estimation [21]; such features are not available in the other open-source libraries. There is ongoing work in the development of an associated graphical user interface extension in 3D slicer for SPECT reconstruction [22]; the corresponding repository can be found on the PyTomography Github organization page.
@@
Overall, the aim of this study is to introduce the community to the PyTomography software project, demonstrate its current capabilities and flexibility, and to encourage community involvement in the continuation of its development. Example applications are explored in SPECT and PET reconstruction, though the library also has capabilities for clinical CT reconstruction as well. In what follows, we elaborate on our methods and results for PyTomography.
@@ 2.1 Software scope and functionality
A brief overview of the mathematical paradigm of medical imaging is presented in order to motivate the software architecture. The mathematical conventions used in this paper are as follows. Vectors are represented by lower case letters, while linear operators are represented by upper case letters. The product of two vectors $vw$ and division of two vectors $v/w$ are inferred to be point-wise operations. Components of a vector are given by $v_{i}$, and components of linear operators are given by $A_{ij}$. When $A1$ or $A^{T}1$ is written, “1” is to be interpreted as a vector (in the appropriate vector space) with all elements equal to 1. When classes/functions are quoted, they refer to the either (i) functionality of the current PyTomography version or (ii) functionality of dependencies (see pyproject.toml file) of the current PyTomography version.
@@ 2.1.1 Projectors
Linear imaging systems can be characterized by a system matrix $H:\mathcal{U}\to\mathcal{V}$. Its components $H_{ij}$ represent the contribution of element $j$ in object space to element $i$ in projection space. In emission tomography, analytical modeling of the data expectation $\bar{g}$ can be achieved via $\bar{g}=Hf+s$ where $s$ is an additive term containing estimates for scattered events (SPECT) or scatter and random coincidences (PET). The operation of $H$ on an object vector is known as forward projection (FP), while the operation of $H^{T}$ on a projection vector is known as back projection (BP). Example implementations available in PyTomography are listed below:
@@
    where $n$ designates the projection angle, $\mathcal{T}_{\text{rot}}^{(n)}$ implements a rotation, $\mathcal{T}_{\text{trans}}^{(n)}$ implements a translation that accounts for non-central projections, $\mathcal{T}_{\text{att}}^{(n)}$ implements the adjustment for photon attenuation, $\mathcal{T}_{\text{cdr}}^{(n)}$ implements depth-dependent collimator blurring, $\mathcal{P}_{x}$ is summation along the $+x$ axis, $\mathcal{T}_{\text{fov}}^{(n)}$ applies a mask in projection space that accounts for a finite field of view, $v_{n}$ is a unit vector of length $N$ with 1 in component $n$, and $\otimes$ is the outer product. The most computationally expensive operations in this sequence are $\mathcal{T}_{\text{cdr}}^{(n)}$ and $\mathcal{T}_{\text{rot}}^{(n)}$, which are implemented using the PyTorch functions rotate (from torchvision.transforms.functional) and conv1d/conv2d (from torch.nn).
@@
    where $\mathcal{T}_{rr}$ implements object-space resolution modeling (typically via a uniform Gaussian kernel), $\mathcal{P}_{\Sigma}$ implements line integral computation (listmode/sinogram), $\mathcal{T}_{\eta}$ adjusts for detector normalization, and $\mathcal{T}_{\omega}$ adjusts for attenuation effects. The most computationally expensive operation is $\mathcal{P}_{\Sigma}$, which is implemented using joseph3d_fwd (non-TOF) and joseph3d_fwd_tof_sino/ joseph3d_fwd_tof_lm (TOF) from parallelproj.
PyTomography can be used to construct projectors that extend projection space using the ExtendedSystemMatrix class; such functionality is necessary, for example, when modeling multiple photopeaks in SPECT or when using motion correction. In such situations, the extended projection space has dimensionality $(N,...)$ where ... is the original dimensionality of projection space, and $N$ represents the number of photopeaks or time gates during acquisition. Mathematically, these projectors can be written as
@@
    Multiple SPECT Photopeaks: In this scenario, $A_{n}$ are scaling factors that adjust for detector sensitivities and branching ratios between different emission peaks, $B_{n}$ are identity mappings, and $H_{n}$ represents the system matrix corresponding to a particular photopeak. The extended projection space has dimensions $(N,...)$ where $N$ is the number of different photopeaks.
@@ 2.1.2 Likelihoods
A statistical likelihood function $\tilde{L}(g|\bar{g})$ is proportional to the probability of observing an actual realization of data $g$ given expectation $\bar{g}$. Since $\bar{g}$ is approximated using $f$ and $H$, it follows that $\tilde{L}(g|\bar{g})=\tilde{L}(g|f,H,...)$, where the ellipsis “...” contains additional parameters such as a scatter estimate $s$; the ellipsis will be assumed implicit for the remainder of this paper. Many iterative reconstruction algorithms seek to maximize the likelihood function via gradient ascent by exploiting the gradient $\nabla_{f}\tilde{L}(g|f,H)$. Since (i) computation of $\nabla_{f}\log\tilde{L}(g|f,H)$ is more tenable than $\nabla_{f}\tilde{L}(g|f,H)$ and (ii) maximization of $\log\tilde{L}(g|f,H)$ is equivalent to maximization of $\tilde{L}(g|f,H)$, the statistical “log-likelihood” is nearly always considered instead. In PyTomography, the “likelihood” $L(g|f,H)$ refers to an objective function to be maximized; as such, the “likelihood” $L(g|f,H)$ is often used interchangeably with the statistical “log-likelihood” $\log\tilde{L}(g|f,H)$ for a particular modality. Likelihood functions for different modalities are shown in Table 2; evaluation of a likelihood and its derivatives require forward and back projection of system matrices.
@@ 2.1.3 Reconstruction Algorithms
where the ellipsis “...” includes all additional hyperparameters required for the algorithm; the ellipsis will be assumed implicit for the remainder of this paper. The main algorithms in PyTomography are preconditioned gradient ascent (PGA) algorithms. These algorithms aim to maximimize the likelihood via iterative updates that take the form
@@
where $n$ is the subiteration index (used as a superscript and not an exponent for all variables), $f^{n}$ is the object estimate of subiteration $n$, $g^{n}$ specifies a subset of data used in subiteration $n$, $H^{n}$ is a system matrix projecting only to the current subset, $C^{n}(f^{n})$ is a preconditioner factor that depends on the object estimate, $V(f^{n})$ is a prior function used to regularize the algorithm, and $\beta^{n}=c^{n}\beta$ is a scaling factor for the prior function (scaled by the fraction of subset data $c^{n}$). These algorithms partition the data $g$ into $M$ approximately equal subsets; a single iteration of an algorithm is defined as the number of subiterations where all subsets have been used once. Various PGA algorithms available in PyTomography are shown in Table 2; they are characterized by their preconditioner factor.
@@
where $w_{rs}$ provides a weighting factor between the locations voxel $r$ and voxel $s$ and $\phi(f_{r},f_{s})$ characterizes a weighting factor between the observed image values at voxel $r$ and voxel $s$. When $w_{rs}$ is non-zero only for neighbouring voxels, the prior is known as a nearest neighbour prior; various nearest neighbour priors implemented in PyTomography are shown in Table 2. While the weighting factor $w_{rs}$ is typically proportional to inverse Euclidean distance, it may also be dependent on an additional anatomical image $\chi$ whereby $w_{rs}=w_{rs}(\chi_{r},\chi_{s})$ Since anatomical information is used, this is commonly referred to as an anatomical prior (AP); variations of this technique have recently become popular in clinical practice due to enhanced lesion quantitation and detectability in bone SPECT/CT [6]. In this paper, use of an AP refers to using only the 8 most similar neighbours based on absolute differences of HU values in a CT scan or attenuation map. The prior used in subsequent examples in this paper is the relative difference penalty (RDP) [23], since it is commonly used in vendor reconstruction.
Table 2: Sample PyTomography classes and their corresponding identifying features. $s$ corresponds to additive term in likelihoods. The sample algorithms shown are the ordered subset expectation maximum (OSEM) [24], ordered subset maximum a posteriori one step late (OSMAPOSL) [25], block sequential regularized expectation maximum (BSREM) [26]. $n$ is used as a superscript (not an exponent) for all variables. For BSREM, $\alpha^{n}$ is a relaxation parameter. The sample nearest neighbour prior functions are characterized by their image weighting function $\phi(f_{r},f_{s})$.
Projector classes System matrix $H$
SPECTSystemMatrix Equation 1
PETLMSystemMatrix Equation 2
Likelihood classes Gradient $\nabla_{f}L(g|f,H)$
PoissonLogLikelihood $H^{T}\left(\frac{g}{Hf+s}\right)-H^{T}1$
NegativeMSELikelihood $H^{T}W(g-(Hf+s))$
Algorithm classes Preconditioner factor $C^{n}(f^{n},H^{n})$
OSEM $f^{n}/\left(\left(H^{n}\right)^{T}1\right)$ and $V(f^{n})=0$
OSMAPOSL $f^{n}/\left(\left(H^{n}\right)^{T}1+\nabla_{f}V(f^{n})\right)$
BSREM $f^{n}\alpha^{n}/\left(c^{n}H^{T}1\right)$
Prior classes Weighting function $\phi(f_{r},f_{s})$
QuadraticPrior $\left((f_{r}-f_{s})/\delta\right)^{2}/4$
LogCoshPrior $\tanh((f_{r}-f_{s})/\delta)$
RelativeDifferencePrior $\left((f_{r}-f_{s})^{2}\right)/\left(f_{r}+f_{s}+\gamma|f_{r}-f_{s}|\right)$
@@ 2.2 Software architecture
    Data/Metadata. The standard data type used to represent images and raw projection data is the tensor class of PyTorch; data can be stored on CPU or GPU. Metadata classes contain all auxiliary information needed for image reconstruction pertaining to a particular imaging system, such as voxel size, spatial dimensions, and information pertaining to the raw data of a particular modality. The input/output (I/O) module is used for reading various file formats; it currently supports SIMIND, DICOM (SPECT/CT), GATE (PET), and HDF5 listmode (PET) formats. Data reading consists of loading the generated files and constructing the appropriate data/metadata instances within the PyTomography framework. The I/O module is also used to read internal files of the library, such as coefficients to convert Hounsfield Units to attenuation coefficients, files containing SPECT collimator information, and files containing PET scanner geometry.
@@
    Projectors. Projectors are mathematical operators that map between object and projection space. The system matrix is an example of a projector that models a particular imaging system. While the input/output datatype to projectors is the PyTorch tensor class, internal code can utilize functionality from other libraries. Projector subclasses inherit from a generic SystemMatrix class and implement a forward, backward, and compute_normalization_factor method. Projectors used in subset based algorithms should additionally implement set_n_subsets, get_projection_subset and get_weighting_subset; the second method takes in projection data and returns subset data based on how the projector defines subsets, and the third gets the corresponding weight ($c^{n}$) for scaling preconditioners.
@@
    Likelihoods. Likelihoods specify the likelihood function that characterizes the acquired data. Likelihoods $L(g|H,f)$ are initialized using a system matrix $H$, projection data $g$, and any additional dependencies; once initialized they act as functions of an image estimate $f$. Likelihoods inherit from a generic Likelihood class and should implement the compute_gradient method. Higher order derivatives of the likelihood may also be defined as separate methods within the class.
@@
    Algorithms. Algorithms are used for image reconstruction, and are designed to interface with all imaging modalities. Statistical iterative algorithms require as input a likelihood instance, and may take as optional arguments prior functions and callbacks. Linear PGAAs inherit from a generic LinearPreconditionedGradientAscentAlgorithm class and simply need to implement the _linear_preconditioner_factor and optionally re-implement the __init__ method to account for any additional required arguments.
Sample code demonstrating the modularity of the library is shown in Figure 1 for a simplified SPECT use case (e.g. no attenuation, PSF, or scatter correction). Lines 1-9 demonstrate reading the data/metadata, lines 10-13 demonstrate constructing the projector using the loaded data/metadata, lines 14-17 demonstrate creating the likelihood using the system matrix and loaded data, and lines 18-22 demonstrate creating a reconstruction algorithm and obtaining a reconstructed image. A new contribution to the library might involve adding a new system model and replacing lines 1-13 while reusing lines 14-22. Another possible contribution might include a new reconstruction algorithm; in this case lines 18-22 would need to be replaced while the SPECT projector code can be reused.
Figure 1: Sample code for a simplified SPECT reconstruction use case (e.g. no attenuation/PSF/scatter correction) that demonstrates the modular structure of the library.